OpenStudy (anonymous):

please help factor this expression. x^2+11x+28

OpenStudy (mathmate):

for x^2+11x+28 a=1 b=11 c=28 We need to find m,n such that m+n=b=11 m*n=ac=28 m=7, n=4 fit the bill. x^2+11x+28=(x+7)(x+4)

OpenStudy (anonymous):

thank you mate! :D

OpenStudy (mathmate):

You're welcome! :)

OpenStudy (anonymous):

one question, do you know any page or something that could have the direct answer? i need to check